OpenStudy (crashonce):

1+lgx=lg(11x-3) find x @ParthKohli

Parth (parthkohli):

\[1 + \log(x) = \log(10) + \log(x) = \log(10x)\]

OpenStudy (kainui):

Assuming base e... \[\LARGE 1+\ln x=\ln(11x-3)\] we can say that \[\LARGE 1=\ln e\] and get \[\LARGE \ln e+\ln x=\ln(11x-3) \\ \LARGE ex=11x-3\]

OpenStudy (kainui):

After posting it seems obvious we're dealing with base 10. Whatever, still the same idea.

OpenStudy (perl):